The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 66 September 30, 2002

Contents

Letter From The Editor  1
Abstracts of 10 Refereed Papers 2
Recent astro-ph Listings  9

Letter From The Editor

Dear Colleagues,

It gives us great pleasure to announce that we now have a fully-functional online submission system, available from our Website www.astro.uiuc.edu/projects/mcnews/MCNews.html. This is the first step in enhancements and upgrades we will be bringing to the Magellanic Clouds Newsletter in the near future. In addition, we are adding a new feature to MC News, a listing of recent submissions to astro-ph dealing with the Magellanic Clouds. We hope this new feature will be useful to our readers. We have not included full abstracts of the astro-ph listings, nor is our search-method 100% successful in selecting Magellanic Cloud-related materials, so we do encourage readers of MC News to continue to submit abstracts.

Best Wishes, Bryan Dunne MC News Editor

Abstracts of Refereed Papers

Recombination-line vs. forbidden-line abundances in H ii regions

Yiannis G. Tsamis^{1}, M. J. Barlow^{1}, X.-W. Liu^{1}, I. J. Danziger^{2} and P. J. Storey^{1}

1Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 2Osservatorio Astronomico di Trieste, Via G. B. Tiepolo 11, I-34131 Trieste, Italy

We have obtained deep optical, long-slit spectrophotometry of the Galactic H ii regions M 17, NGC 3576 and of the Magellanic Cloud H ii regions 30 Doradus, LMC N11B and SMC N66, recording the optical recombination lines (ORLs) of C ii, N ii and O ii. A spatial analysis of 30 Doradus is performed, revealing that the forbidden-line [O iii] electron temperature is remarkably constant across the nebula. The forbidden-line O^{2+}/H^{+} abundance mapped by the [O iii] λ4959 collisionally excited line (CEL) is shown to be consistently lower than the recombination-line abundance mapped by the O ii V 1 multiplet at 4650 Å. In addition, the spatial profile of the C^{2+}/O^{2+} ratio derived purely from recombination lines is presented for the first time for an extragalactic nebula. Temperature-insensitive ORL C^{2+}/O^{2+} and N^{2+}/O^{2+} ratios are obtained for all nebulae except SMC N66. The ORL C^{2+}/O^{2+} ratios show remarkable agreement within each galactic system, while also being in agreement with the corresponding CEL ratios. The disagreement found between the ORL and CEL N^{2+}/O^{2+} ratios for M 17 and NGC 3576 can be attributed to the N ii V 3 and V 5 ORLs that were used being affected by fluorescent excitation effects. For all five nebulae, the O^{2+}/H^{+} abundance derived from multiple O ii ORLs is found to be higher than the corresponding value derived from the strong [O iii] λλ4959, 5007 CELs, by factors of 1.8–2.7 for four of the nebulae. The LMC N11B nebula exhibits a more extreme discrepancy factor for the O^{2+} ion, ∼5. Thus these H ii regions exhibit ORL/CEL abundance discrepancy factors that are similar to those previously encountered amongst planetary nebulae. Our optical CEL O^{2+}/H^{+} abundances agree to within 20-30 per cent with published O^{2+}/H^{+} abundances that were obtained from observations of infrared fine-structure lines. Since the low excitation energies of the latter make them insensitive to variations about typical nebular temperatures, fluctuations in temperature are ruled out as the cause of the observed ORL/CEL O^{2+} abundance discrepancies. We present evidence that the observed O ii ORLs from these H ii regions originate from gas of very similar density (< 3500 cm^{−3}) to that emitting the observed heavy-element optical and infrared CELs, ruling out models that employ high-density ionized inclusions in order to explain the abundance discrepancy. We consider a scenario whereby much of the heavy-element ORL emission originates from cold (≤ 500 K) metal-rich ionized regions. These might constitute halos that are being evaporated from much denser neutral cores. The origin of these metal-rich inclusions is not clear – they may have been ejected into the nebula by evolved, massive Of and Wolf-Rayet stars, although the agreement found between heavy element ion ratios derived from ORLs with the ratios derived from CELs provides no evidence for nuclear-processed material in the ORL-emitting regions.

Comments: Accepted by MNRAS WWW: http://arXiv.org/abs/astro-ph/0209534 e-mail: mjb@star.ucl.ac.uk

Wolf-Rayet Binaries in the Magellanic Clouds and Implications for Massive-Star Evolution I. SMC

C. Foellmi (1 and 2), A.F.J. Moffat (1), M.A. Guerrero (3)

(1) Université de Montréal et Observatoire du Mont Mégantic, C.P. 6128, Succ. Centre-Ville, Montréal (Qc) H3C 3J7, Canada

(2) European Southern Observatory, Alonso de Cordova 3107, casilla 19001, Vitacura, Santiago 19, Chile

(3) Department of Astronomy, University of Illinois at Urbana-Champaign, 1002 W. Green St., Urbana, IL 61801, USA

We have carried out an intensive spectroscopic campaign to search for binaries via periodic radialvelocity (RV) variations among all the nitrogen-rich WN Wolf-Rayet stars in the Small Magellanic Clouds (SMC), and all WNE stars in the Large Magellanic Cloud (LMC). We present in this first paper the results for the SMC. Along with the results of Bartzakos, Moffat & Niemela (2001) on the only carbon/oxygen-rich WR star (AB8, WO4+O4), the whole WR population of the SMC (11 stars) has now been investigated intensively for periodic RV variability. We have also retrieved time-dependent photometric data in the public domain from the OGLE and MACHO projects, as well as X-ray data from the ROSAT and Chandra archives, to provide additional constraints on the binary character. Contrary to theoretical expectations which predict a virtually 100% binary frequency in the SMC, we find a normal (∼ 40%) WR binary frequency in this galaxy. We also find the clear presence of hydrogen in the winds of the single WR stars in the SMC, even for the stars with an early spectral subtype. We discuss the possible reasons and implications of this for stellar evolution of massive stars in such a low-metallicity environment, e.g. the influence of rotation versus the necessity of very high initial mass of the progenitors for single stars, and the possible past occurrence of Roche-lobe overflow for binaries.

Comments: Accepted by MNRAS Contact: cfoellmi@eso.org WWW: http://www.astro.umontreal.ca/ foellmi/papers/smc.pdf e-mail: cfoellmi@eso.org

Wolf-Rayet Binaries in the Magellanic Clouds and Implications for Massive-Star Evolution II. LMC

C. Foellmi (1 and 2), A.F.J. Moffat (1), M.A. Guerrero (3)

(1) Université de Montréal et Observatoire du Mont Mégantic, C.P. 6128, Succ. Centre-Ville, Montréal (Qc) H3C 3J7, Canada

(2) European Southern Observatory, Alonso de Cordova 3107, casilla 19001, Vitacura, Santiago 19, Chile

(3) Department of Astronomy, University of Illinois at Urbana-Champaign, 1002 W. Green St., Urbana, IL 61801, USA

We present in this second paper the results for the Large Magellanic Cloud (LMC) of our intensive spectroscopic campaign to search for binaries via periodic radial-velocity (RV) variations among Wolf-Rayet (WR) stars. We observed 61 nitrogen-rich WNE stars in the LMC. Along with the results of Bartzakos et al. (2001) on the carbon/oxygen-rich WR stars, 2/3 of the WR population of the LMC (134 stars in total) has now been investigated for periodic RV variability. We have also retrieved time-dependent photometric data in the public domain from the OGLE and MACHO projects, as well as X-ray data from ROSAT and Chandra satellites, to provide additional constraints on the binary character. For each of our sample stars, we discuss its observational properties: RV variations, (periodic) photometric variability, X-ray luminosity, spectral classification, abundance of hydrogen, runaway status and line-profile variations (LPVs). For the binaries we discuss additional properties, like wind-wind collision (WWC) effects, and the orbital parameters. With this large sample, we discuss the global properties of the WNE population, which is expected to be the most sensitive to binary evolution with respect to the influence of metallicity. To emphasize the relevance of the binary frequency test for the stellar evolution of massive stars in the LMC, we review their observational properties and provide new and meaningful evolutionary classes, which reconcile observational and theoretical definitions. Finally, we draw an overall evolutionary scheme for massive-star evolution, with respect to the three main ingredients of stellar evolution: mass, metallicity and rotation.

Comments: Submitted to MNRAS Contact: cfoellmi@eso.org WWW: http://www.astro.umontreal.ca/ foellmi/papers/lmc.pdf e-mail: cfoellmi@eso.org

An X-ray investigation of the NGC346 field in the SMC (1) : the LBV HD5980 and the NGC346 cluster

Y. Nazé^{1}, J.M. Hartwell^{2}, I.R. Stevens^{2}, M. F. Corcoran^{3},Y.-H. Chu^{4}, G. Koenigsberger^{5}, A.F.J. Moffat^{6}, and V.S. Niemela^{7}

1 Institut d’Astrophysique et de Géophysique, Université de Liège, Allée du 6 Août 17, Bat. B5c, B 4000 - Liège (Belgium) 2 School of Physics & Astronomy, University of Birmingham, Edgbaston, Birmingham B15 2TT (UK) 3 Universities Space Research Association, High Energy Astrophysics Science Archive Research Center, Goddard Space Flight Center, Greenbelt, MD 20771 (USA) 4 Astronomy Department, University of Illinois, 1002 W. Green Street, Urbana, IL 61801 (USA) 5 Instituto de Astronomía, Universidad Nacional Autónoma de Mexico, Apdo. Postal 70-264, 04510, Mexico D.F. (Mexico) 6 Département de physique, Université de Montreal, C.P. 6128, Succ. Centre-Ville, Montreal, QC, H3C 3J7 (Canada) 7 Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de la Plata, Paseo del Bosque S/N, B19000FWA La Plata (Argentina)

We present results from a Chandra observation of the NGC346 cluster. This cluster contains numerous massive stars and is responsible for the ionization of N66, the most luminous HII region and the largest star formation region in the SMC. In this first paper, we will focus on the characteristics of the main objects of the field. The NGC346 cluster itself shows only relatively faint X-ray emission (with L_{X}^{unabs} ∼ 1.5 × 10^{34} erg s^{−1}), tightly correlated with the core of the cluster. In the field also lies HD5980, a LBV star in a binary (or possibly a triple system) that is detected for the first time at X-ray energies. The star is X-ray bright, with an unabsorbed luminosity of L_{X}^{unabs} ∼ 1.7 × 10^{34} erg s^{−1}, but needs to be monitored further to investigate its X-ray variability over a complete 19d orbital cycle. The high X-ray luminosity may be associated either with colliding winds in the binary system or with the 1994 eruption. HD5980 is surrounded by a region of diffuse X-ray emission, which is a supernova remnant. While it may be only a chance alignment with HD5980, such a spatial coincidence may indicate that the remnant is indeed related to this peculiar massive star.

Comments: Accepted by: ApJ WWW: http://www.arXiv.org/abs/astro-ph/0208289 e-mail: naze@astro.ulg.ac.be

Structure and Dynamics of Candidate O Star Bubbles in N44

Yaël Nazé^{1,5}, You-Hua Chu^{2,6}, Martín A. Guerrero^{2,6}, M. S. Oey^{3}, Robert A. Gruendl^{2}, R. Chris Smith^{4}

1 Institut d’Astrophysique et de Géophysique, Université de Liège, Allée du 6 Août 17, Bat. B5c, B 4000 -Liège, Belgium 2 Astronomy Department, University of Illinois, 1002 W. Green Street, Urbana, IL 61801, USA 3 Lowell Observatory, 1400 West Mars Hill Rd., Flagstaff, AZ 86001, USA 4 Cerro Tololo Inter-American Observatory, Casilla 603, La Serena, Chile 5 Research Fellow FNRS (Belgium) 6 Visiting astronomer, Cerro Tololo Inter-American Observatory

Dynamical studies of superbubbles and Wolf-Rayet ring nebulae show discrepancies from the standard, adiabatic model for wind-blown bubbles. We therefore study the physical properties and kinematics of three candidate bubbles blown by single O stars, to evaluate whether these discrepancies are also found in these simpler objects. Our sample candidates are N44F, N44J, and N44M, in the outskirts of the HII complex N44 in the Large Magellanic Cloud. We have obtained ground-based and HST emission-line images and high dispersion echelle spectra for these objects. From the Hα luminosities and the [OIII]/Hα ratios of these nebulae, we estimate the spectral types of the ionizing stars to be O7V, O9.5V and O9.5V for N44F, N44J, and N44M, respectively. We find that the observed expansion velocity of 12 km/s for N44F is consistent with the stellar wind luminosity expected from the central ionizing star, as predicted by the standard bubble model. The observed upper limits for the expansion velocities of N44J and N44M are also compatible with the expected values, within the uncertainties. We also report the discovery in N44F of strongly-defined dust columns, similar to those seen in the Eagle Nebula. The photoevaporation of these dense dust features may be kinematically important and may actually govern the evolution of the shell. The inclusion of photoevaporation processes may thus undermine the apparent agreement between the observed bubble dynamics and the simple adiabatic models.

Comments: Accepted by: AJ WWW: http://www.arXiv.org/abs/astro-ph/0208430 e-mail: naze@astro.ulg.ac.be

Optical counterparts to four X-ray sources in the Small Magellanic Cloud

W. R. T. Edge, M. J. Coe

Department of Physics and Astronomy, Southampton University

We report on the spectroscopic and photometric analysis of possible optical counterparts to four X-ray sources in the Small Magellanic Cloud: AX J0049.4-7323, AX J0057.4-7325, RX J0058.2-7321 and RX J0101.1-7206. In the case of the last two, we suggest that the presence of strong Halpha emission from previously proposed optical candidates is definite confirmation that these are Be stars. Similarly, we detected strong Halpha emission from the optical source identified with RX J0049.7-7323 within the error circle for AX J0049.4-7323 and we conclude that these are one and the same object. We were unable to detect any Halpha emission from any of the candidates for AX J0057.4-7325 and the associated photometric analysis was also inconclusive.

Comments: This paper has been accepted by MNRAS. WWW: http://xxx.soton.ac.uk/abs/astro-ph/0209327 e-mail: wrte@astro.soton.ac.uk

Pulsation at the tip of the first giant branch?

Yoshifusa Ita^{1}, Toshihiko Tanabé^{1}, Noriyuki Matsunaga^{1}, Yasushi Nakajima^{2}, Chie Nagashima^{2}, Takahiro Nagayama^{2}, Daisuke Kato^{2}, Mikio Kurita^{2}, Tetsuya Nagata^{2},

Shuji Sato^{2}, Motohide Tamura^{3}, Hidehiko Nakaya^{4} and Yoshikazu Nakada^{1,5}

1Institute of Astronomy, School of Science, The University of Tokyo, Mitaka, Tokyo 181-0015, Japan 2Department of Astrophysics, Nagoya University, Chikusa-ku, Nagoya 464-8602, Japan 3National Astronomical Observatory of Japan, Mitaka, Tokyo 181-8588, Japan 4Subaru Telescope, National Astronomical Observatory of Japan, 650 North A’ohoku Place, Hilo, HI 96720, U.S.A. 5Kiso Observatory, School of Science, The University of Tokyo, Mitake, Kiso, Nagano 397-0101, Japan

The first results of our ongoing near-infrared (NIR) survey of the variable red giants in the Large Magellanic Cloud, using the Infrared Survey Facility (IRSF) and the SIRIUS infrared camera, are presented. In the observed colour-magnitude diagram (J − K, K) and the stellar K magnitude distribution, the tip of the first giant branch (TRGB), where is the starting point of helium burning in the core is clearly seen. Apart from the genuine AGB variables, we found many variable stars that populate just around the TRGB. From these results, we infer that most, if not all, of them are RGB variables.

Comments: Submitted to: MNRAS WWW: http://www.ioa.s.u-tokyo.ac.jp/ yita/scr/mn/yita2.ps.gz e-mail: yita@ioa.s.u-tokyo.ac.jp

Surface brightness profiles and structural parameters for 53 rich stellar clusters in the Large Magellanic Cloud

A. D. Mackey and G. F. Gilmore

Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK

We have compiled a pseudo-snapshot data set of two-colour observations from the Hubble Space Telescope archive for a sample of 53 rich LMC clusters with ages 10^{6} − 10^{10} yr. We present surface brightness profiles for the entire sample, and derive structural parameters for each cluster, including core radii, and luminosity and mass estimates. Because we expect the results presented here to form the basis for several further projects, we describe in detail the data reduction and surface brightness profile construction processes, and compare our results with those of previous ground-based studies. The surface brightness profiles show a large amount of detail, including irregularities in the profiles of young clusters (such as bumps, dips, and sharp shoulders), and evidence for both double clusters and post core-collapse (PCC) clusters. In particular we find power-law profiles in the inner regions of several candidate PCC clusters, with slopes of approximately −0.7, but showing considerable variation. We estimate that 20 ± 7 per cent of the old cluster population of the LMC has entered PCC evolution, a similar fraction to that for the Galactic globular cluster system. In addition, we examine the profile of R136 in detail and show that it is probably not a PCC cluster. We also observe a trend in core radius with age that has been discovered and discussed in several previous publications by different authors. Our diagram has better resolution however, and appears to show a bifurcation at several hundred Myr. We argue that this observed relationship reflects true physical evolution in LMC clusters, with some experiencing small scale core expansion due to mass loss, and others large scale expansion due to some unidentified characteristic or physical process.

Comments: Accepted by MNRAS, and is astro-ph/0209031 WWW: http://www.ast.cam.ac.uk/STELLARPOPS/LMC clusters/ e-mail: dmackey@ast.cam.ac.uk

Surface brightness profiles and structural parameters for 10 rich stellar clusters in the Small Magellanic Cloud

A. D. Mackey and G. F. Gilmore

Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK

As a follow up to our recent study of a large sample of LMC clusters (Mackey & Gilmore 2002a) , we have conducted a similar study of the structures of ten SMC clusters, using archival Hubble Space Telescope snapshot data. We present surface brightness profiles for each cluster, and derive structural parameters, including core radii and luminosity and mass estimates, using exactly the same procedure as for the LMC sample. Because of the small sample size, the SMC results are not as detailed as for the larger LMC sample. We do not observe any post core-collapse clusters (although we did not expect to), and there is little evidence for any double clusters in our sample. Nonetheless, despite the small sample size, we show for the first time that the SMC clusters follow almost exactly the trend in core radius with age observed for the LMC system, including the apparent bifurcation at several hundred Myr. This further strengthens our argument that this relationship represents true physical evolution in these clusters, with some developing significantly expanded cores due to an as yet unidentified physical process. Additional data, both observational and from N-body simulations, is still required to clarify many issues.

Comments: Accepted to MNRAS, is also astro-ph/0902046 WWW: http://www.ast.cam.ac.uk/STELLARPOPS/SMC clusters/ e-mail: dmackey@ast.cam.ac.uk

The Master Catalogue of stars towards the Magellanic Clouds*

N. Delmotte^{1,4}, C. Loup^{2}, D. Egret^{1}, M.-R. Cioni^{3,4}, F. Pierfederici^{4}

1 CDS, Observatoire Astronomique de Strasbourg, UMR 7550, Université Louis Pasteur, 67000 Strasbourg, France 2 Institut d’Astrophysique de Paris, CNRS UPR 341, 98 bis Bd Arago, 75014 Paris, France 3 Leiden Observatory, University of Leiden, P.O. Box 9513, 2300 RA Leiden, The Netherlands 4 European Southern Observatory, ESO, K.-Schwarzschild-Str.-2, D-85748 Garching, Germany

The Master Catalogue of stars towards the Magellanic Clouds (MC2)^{∗∗} is a multi–wavelength reference catalogue. The current paper presents the first results of the MC2 project. We started with a massive cross–identification of the two recently released near–infrared surveys: the DENIS Catalogue towards the Magellanic Clouds (DCMC) with more than 1.3 million sources identified in at least two of the three DENIS filters (I J Ks) and the 2nd Incremental Release of the 2MASS point source catalogue (J H Ks) covering the same region of the sky. Both point source catalogues provide an unprecedented wealth of data on the stellar populations of the Magellanic Clouds (MCs). The cross– matching procedure has be en extended to optical wavelength ranges, including the UCAC1 (USNO) and GSC2.2 catalogues. New cross–matching procedures for very large catalogues have been developed and important results o n the astrometric and photometric accuracy of the cross–identified catalogues were derived. The cross–matching o f large surveys is an essential tool to improve our understanding of their specific contents. This study has been partly supported by the astrovirtel^{∗∗∗} project that aims at improving access to astronomical archives as virtual telescopes.

∗∗http : //vizier.u − strasbg.fr/MC2/
∗∗∗http : //www.stecf.org/astrovirtel/

Comments: *This work has been supported by astrovirtel, a project funded by the European Commission under FP5 Contract No. HPRI-CT-1999-00081. Accepted by: Astronomy & Astrophysics WWW: http://vizier.u-strasbg.fr/MC2/ e-mail: delmotte@astro.u-strasbg.fr

Recent astro-ph Listings

1. astro-ph/0209626:
Title: NGC1866: New analysis from HST-WFPC2 observations of the inner regions
Authors: E. Di Carlo (1), G. Raimondo (1), E. Brocato (1), V. Castellani (2), A.R. Walker (3)
((1) INAF-Osservatorio Astronomico di Collurania, Teramo, Italy; (2) INAF-Osservatorio
Astronomico di Monte Porzio, Roma, Italy; (3) CTIO-American Observatory, NOAO, La
Serena, Chile)
Comments: 2 pages, incl. 2 figures, uses newpasp.sty, to be published in ``New Horizons in
Globular Cluster Astronomy'', ASP Conference Series, 2002; Eds.: G. Piotto, G. Meylan, G.
Djorgowski and M. Riello
2. astro-ph/0209580:
Title: Modelling the Recurrent Nova CI Aql in Quiescence
Authors: C. Lederle, S. Kimeswenger (University Innsbruck)
Comments: 9 pages 14 eps figures, to appear in Astron. & Astrophys
3. astro-ph/0209534:
Title: Heavy elements in Galactic and Magellanic Cloud HII regions: recombination-line
versus forbidden-line abundances
Authors: Yiannis G. Tsamis (1), M. J. Barlow (1), X.-W. Liu (1), I. J. Danziger (2), P. J.
Storey (1) (1 Department of Physics and Astronomy, University College London, UK) (2
Osservatorio Astronomico di Trieste, Italy)
Comments: 24 pages; 9 figures; accepted by Monthly Notices of the Royal Astronomical
Society
4. astro-ph/0209523:
Title: Detection of Carbon Monoxide within the Magellanic Bridge
Authors: E. Muller, L. Staveley-Smith, W. Zealey
Comments: 8 pages, 6 Figures. LaTeX. Accepted for publication by MNRAS
5. astro-ph/0209513:
Title: Variability-Selected Quasars in MACHO Project Magellanic Cloud Fields
Authors: M. Geha, C. Alcock, R.A. Allsman, D.R. Alves, T.S. Axelrod, A.C. Becker, D.P.
Bennett, K.H. Cook, A.J. Drake, K.C. Freeman, K. Griest, S.C. Keller, M.J. Lehner, S.L.
Marshall, D. Minniti, C.A. Nelson, B.A. Peterson, P. Popowski, M.R. Pratt, P.J. Quinn, C.W.
Stubbs, W. Sutherland, A.B. Tomaney, T. Vandehei, D.L. Welch (The MACHO
Collaboration)
Comments: 26 pages, 9 figures. High resolution figures and data available from this http
URL Accepted for publication in the Astronomical Journal
6. astro-ph/0209481:
Title: Synthetic High-Resolution Line Spectra of Star-Forming Galaxies Below 1200A
Authors: Carmelle Robert, Anne Pellerin, Alessandra Aloisi, Claus Leitherer, Charles
Hoopes, Timothy M. Heckman
Comments: 33 pages including 13 figures, accepted for publication in ApJ
7. astro-ph/0209414:
Title: SMC X-1 As An Intermediate-Stage Flaring X-ray Pulsar
Authors: Dae-Sik Moon, Stephen S. Eikenberry, Ira M. Wasserman (Cornell University)
Comments: 14 pages (5 figures), To appear in ApJ Letters
8. astro-ph/0209405:
Title: Inferring physical conditions in interstellar clouds of H_2
Authors: Matthew K. Browning, Jason Tumlinson, J. Michael Shull (Colorado)
Comments: 33 pages (aastex, manuscript), 9 figures (3 color). accepted to ApJ

9. astro-ph/0209370:
Title: Revealing New Physical Structures in the Supernova Remnant N63A through
Chandra Imaging Spectroscopy
Authors: Jessica S. Warren (1), John P. Hughes (1), Patrick O. Slane (2) ((1) Rutgers
University, (2) Harvard-Smithsonian Center for Astrophysics)
Comments: 18 pages, 5 figures (2 color), accepted for publication in ApJ
10. astro-ph/0209352:
Title: Direct Lens Imaging of Galactic Bulge Microlensing Events
Authors: Cheongho Han (Chungbuk National University, Korea) Heon-Young Chang
(Korea Institute for Advanced Study)
Comments: total 12 pages, including 4 figures, MNRAS, accepted
11. astro-ph/0209327:
Title: Optical counterparts to four X-ray sources in the Small Magellanic Cloud
Authors: W. R. T. Edge, M. J. Coe (Department of Physics and Astronomy, Southampton
University)
Comments: Paper has been accepted by MNRAS. Amended to include a reference
12. astro-ph/0209259:
Title: A large Wolf-Rayet population in NGC300 uncovered by VLT-FORS2
Authors: H.Schild (Zurich), P.A.Crowther (UCL), J.B.Abbott (UCL), W. Schmutz (Davos)
Comments: 12 pages, 12 figures, submitted to A&A (includes aa.cls) - version with higher
resolution finding charts available from this ftp URL
13. astro-ph/0209255:
Title: Orbital and Super-Orbital Periods of 1E 1740.7-2942 and GRS 1758-258
Authors: D. M. Smith, W. A. Heindl, J. H. Swank
Comments: 13 pages, 5 figures, accepted by The Astrophysical Journal Letters
14. astro-ph/0209207:
Title: The Master Catalogue of stars towards the Magellanic Clouds
Authors: N. Delmotte (1 and 4), C. Loup (2), D. Egret (1), M.-R. Cioni (3 and 4), F.
Pierfederici (4) ((1) CDS, France, (2) IAP, France, (3) Leiden, The Netherlands, (4) ESO)
Comments: 14 pages, 29 figures. Accepted for publication in A&A. For more information,
see this http URL
15. astro-ph/0209193:
Title: The Chemical Uniformity of High z DLA Protogalaxies
Authors: Jason X. Prochaska (UCO/Lick Observatory)
Comments: 13 pages, 5 figures. Uses onecolfloat.sty. Accepted to ApJ, September 3, 2002
16. astro-ph/0209187:
Title: Light Curve Analysis of Be Star Candidates in the Small Magellanic Cloud
Authors: R.E. Mennickent, G. Pietrzynski, W. Gieren
Comments: 3 pages, 7 figures, to be published in ASP Conference Series, Interplay
between periodic, cyclic and stochastic variability in selected areas of the H-R diagram, C.
Sterken (ed.)
17. astro-ph/0209127:
Title: The Magellanic Stream, High-Velocity Clouds and the Sculptor Group
Authors: M.E. Putman (CASA), L. Staveley-Smith (ATNF), K.C. Freeman (RSAA), B.K.
Gibson (Swinburne), D.G. Barnes (Swinburne)
Comments: ApJ accepted, Jan. 2003, for paper with figures included go to this http URL
18. astro-ph/0209060:
Title: Energy Spectra and Normalized Power Spectral Densities of X-Ray Nova GS
2000+25

Authors: K. Terada, S. Kitamoto, H. Negoro, S. Iga
Comments: 24 pages, text only, Full version is published in PASJ 54, 609-627 (2002)
Journal-ref: Publ.Astron.Soc.Jap. 54 (2002) 609-627
19. astro-ph/0209046:
Title: Surface Brightness Profiles and Structural Parameters for 10 Rich Stellar Clusters in
the Small Magellanic Cloud
Authors: A. D. Mackey (1), G. F. Gilmore (1) ((1) IoA, Cambridge, UK)
Comments: Accepted for publication in MNRAS. 11 pages, 4 figures. We recommend that
the interested reader visit this http URL for on-line data and the full resolution paper
20. astro-ph/0209031:
Title: Surface Brightness Profiles and Structural Parameters for 53 Rich Stellar Clusters in
the Large Magellanic Cloud
Authors: A.D. Mackey (1), G.F. Gilmore (1) ((1) IoA, Cambridge, UK)
Comments: Accepted for publication in MNRAS. 24 pages plus 16 figures and 9 pages of
profiles. We strongly recommend that the interested reader visit this http URL for on-line
data and the full resolution paper
21. astro-ph/0208534:
Title: PSR B1951+32 and PSR J0537-6910 - where are the optical counterparts?
Authors: R. F. Butler (National University of Ireland, Galway), A. Golden (National
University of Ireland, Galway), A. Shearer (National University of Ireland, Galway), C.
Gouiffes (CEA/Saclay, France)
Comments: 4 pages, 4 figures, 2 tables. To appear in the Proceedings of the 270.
WE-Heraeus Seminar on Neutron Stars, Pulsars and Supernova Remnants, Jan. 21-25,
2002, Physikzentrum Bad Honnef, eds W.Becker, H.Lesch & J.Truemper. Proceedings are
available as MPE-Report 278
22. astro-ph/0208502:
Title: The Chemical Composition of the 30 Doradus Nebula Derived from VLT Echelle
Spectrophotometry
Authors: Antonio Peimbert
Comments: 37 pages, 2 figures, submitted to the Astrophysical Journal
23. astro-ph/0208471:
Title: AGAPEROS: Searching for variable stars in the LMC Bar II. Temporal and near-IR
analysis of Long-Period Variables
Authors: T. Lebzelter, M. Schultheis, A.L. Melchior
Comments: 12 pages, 15 figures accepted for publication in A&A
24. astro-ph/0208452:
Title: To Circularize or Not to Circularize? -- Orbital Evolution of Satellite Galaxies
Authors: Yoshikazu Hashimoto, Yoko Funato, Junichiro Makino
Comments: AASTEX, 16 pages, 3 figures. Accepted to Astrophisical Journal
25. astro-ph/0208430:
Title: Structure and Dynamics of Candidate O Star Bubbles in N44
Authors: Yael Naze, You-Hua Chu, Martin A. Guerrero, M. S. Oey, Robert A. Gruendl, R.
Chris Smith
Comments: 10 pages, 7 figures (jpg), to be published in AJ. The paper is also available at
this http URL
26. astro-ph/0208417:
Title: Chemical Abundances in Twelve Red Giants of the Large Magellanic Cloud from
High-Resolution Infrared Spectroscopy
Authors: V. V. Smith (U. Texas - El Paso), K. H. Hinkle (NOAO), K. Cunha (Observatorio
Nacional - Brazil), B. Plez (U. Montpellier - France), D. L. Lambert (U. Texas - Austin), C.
A. Pilachowski (Indiana U.), B. Barbuy (U. Sao Paulo - Brazil), J. Melendez (U. Nacional de

San Marcos - Peru), S. Balachandran (U. Maryland), M. S. Bessell (Mt. Stromlo -
Australia), D. P. Geisler (U. Concepcion - Chile), J. E. Hesser (DAO), C. Winge (Gemini
Observatory)
Comments: 25 pages, 6 tables, 10 figures. "In press" in The Astronomical Journal
(scheduled for December 2002)
27. astro-ph/0208399:
Title: Constraining the Location of Microlensing Objects towards the LMC through Parallax
Measurement in EAGLE Observations
Authors: T. Sumi, Y. Kan-ya
Comments: 11 pages, 6 figures. Accepted for publication in MNRAS
28. astro-ph/0208391:
Title: Lithium during the AGB evolution in young open clusters of the Large Magellanic
cloud
Authors: C. Maceroni, V. testa, B. Plez, P. Garcia Lario, F. D'Antona
Comments: 10 pages, 6 figures
29. astro-ph/0208359:
Title: The relative ages of LMC old clusters, and the case of NGC 1841
Authors: Ivo Saviane (1), Alfred Rosenberg (2), Giampaolo Piotto (3), Antonio Aparicio (2)
((1) ESO Chile, (2) IAC, (3) UniPD)
Comments: 2 pages, 1 figure, requires newpasp.sty. To appear in "New Horizons in
Globular Cluster Astronomy", Giampaolo Piotto, George Meylan, George Djorgovski and
Marco Riello, eds
30. astro-ph/0208289:
Title: An X-ray investigation of the NGC346 field in the SMC (1) : the LBV HD5980 and
the NGC346 cluster
Authors: Y. Naze, J.M. Hartwell, I.R. Stevens, M.F. Corcoran, Y.-H. Chu, G.
Koenigsberger, A.F.J. Moffat, V.S. Niemela
Comments: 12 pages, 9 figures (3 in jpg and 6 in ps), to be published in ApJ
31. astro-ph/0208230:
Title: New algorithms and technologies for the un-supervised reduction of Optical/IR
images
Authors: Benoit Vandame
Comments: 12 pages, 9 figures, conference
32. astro-ph/0208188:
Title: Testing intermediate-age stellar evolution models with VLT photometry of LMC
clusters. I. The data
Authors: C. Gallart (IAC), M. Zoccali (ESO), G. Bertelli (Padova), C. Chiosi (Padova), P.
Demarque (Yale), L. Girardi (Trieste), E. Nasi (Padova), J.-H. Woo (Yale), S. Yi (Oxford)
Comments: 27 pages with 10 figures. Accepted by the Astronomical Journal
33. astro-ph/0208162:
Title: The ARAUCARIA project: Deep near-infrared survey of nearby galaxies. I. The
distance to the Large Magellanic Cloud from K-band photometry of red clump stars
Authors: G. Pietrzynski, W. Gieren
Comments: 14 pages, Latex. Accepted for publication in The Astronomical Journal,
November, 2002
34. astro-ph/0208150:
Title: Mass segregation in rich clusters in the Large Magellanic Cloud - III. Implications for
the initial mass function
Authors: R. de Grijs (1), G.F. Gilmore (1), A.D. Mackey (1) M.I. Wilkinson (1), S.F. Beaulieu
(2), R.A. Johnson (3), B.X. Santiago (4) ((1) Cambridge, UK; (2) Victoria, Canada; (3) ESO,
Chile; (4) UFRGS, Brazil)

Comments: 14 pages LaTeX (incl. 9 postscript figures), accepted for publication in MNRAS
35. astro-ph/0208142:
Title: Testing intermediate-age stellar evolution models with VLT photometry of LMC
clusters. II. Analysis with the Yale models
Authors: J.-H. Woo (Yale), C. Gallart (Univ. de Chile), P. Demarque (Yale), S. Yi (Oxford),
M. Zoccali (ESO)
Comments: 19 pages, 13 figures, accepted for the publication in AJ
36. astro-ph/0208057:
Title: Population effects on the red giant clump absolute magnitude: The K-band
Authors: Maurizio Salaris (1), Leo Girardi (2) ((1) Astrophysics Research Institute, Liverpool
John Moores University, UK; (2) INAF - Osservatorio Astronomico di Trieste, Italy)
Comments: 9 pages, 8 figures, MNRAS accepted
37. astro-ph/0208009:
Title: The FUSE survey of OVI absorption in and near the Galaxy
Authors: B.P. Wakker, B.D. Savage (U. Wisconsin), K.R. Sembach (STScI), P. Richter, M.
Meade (U. Wisconsin), E.B. Jenkins (Princeton), J.M. Shull (U.Colorado), T.B. Ake, W.P.
Blair, S.D. Friedman (JHU), J.C. Green (U. Colorado), R.F. Green (NRAO), J.W. Kruk, H.W.
Moos (JHU), E.M. Murphy (U. Virgina), W.R. Oegerle, D.J. Sahnow (JHU), G. Sonneborn
(GSFC), E. Wilkinson (U. COlorado), D.G. York (U. Chicago)
Comments: 85 pages, 127 figures, 13 color figures, 3 tables, AASTeX preprint format. All
figures are in PNG format due to space concerns. Bound copies of manuscript and two
accompanying articles are available upon request. submitted to ApJS